Briefly discuss the price elasticity of demand for tickets from Lusaka to Livingston between 200kwacha and 250 kwacha and clarify how bus fleets revenues are expected to respond to price increases


1
Expert's answer
2021-09-08T18:43:52-0400

"Qd\u200b=2000\u22125P"

"P=200 kwacha"


"Q_d=2000-(5 \times 200)=1000" Q


"Oldpricequantitydemanded=1000"

"If P=250 kwacha"


"=2000\u2212(5\u00d7250)=750"


"New price quantity demanded=1500" Price elasticity of demand Priceelasticityofdemand= "=\frac{\frac{(Q_2-Q_1)}{(Q_1+Q_2)/2}}{\frac{(P_2-P_1)}{(P_2+P_1)/2}}"


="\frac{\frac{(750-1000)}{(750+1000)/2}}{\frac{(250-200)}{(250+200)/2} }"


"=-1.3"


When the bus prices increases, the bus fleets will automatically increase their price. This is because some of the services offered by the fleet may be affected due to increase in price.